A street noodle vendor in Singapore can service an average of 10 customers per hour. Given an average arrival rate of 8 customers per hour, use the Poisson distribution to calculate the probability that the vendor can handle the demand.

What is the probability of having, at most, 10 customers arriving within 1 hour?

a 0.8159
b 0.2834
c 0.1841
d 0.7166
e 0.0993
Answer #1

Answer:

  • Option a: 0.8159

Explanation:

  • For this we will calculate probability of 0, 1, 2, 3, 4, 5,6,7,8,9,10 customers arriving.
  • P(X<=10) = [ P(X=0) + P(X=1) + P(X=2) +P(X=3) + P(X=4) + P(X=5) + P(X=6) + P(X=7) + P(X=8)+ P(X=9)+ P(X=10)]
  • Below is the snapshot of excel for the same (using poisson function in excel)

No of customers Probabilities 0.0003 0.0027 0.0107 0.0286 0.0573 0.0916 0.1221 0.1396 0.1396 0.1241 0.0993 0.8159 0 2 4 5 7 9 10 Sum

Note: Please rate the answer.